Activity   a   m   b   Immediate Predecessor(s)
A   2   6   8   -
B   10   11   12   -
C   1   2   3   -
D   6   7   8   C
E   2   4   6   B, D
F   1   2   3   A, E
G   7   11   13   A, E
H   11   12   13   F
I   5   7   9   G
J   14   16   18   C
K   4   8   12   H, I

a) find expected time and variance
Answer #1

EXPECTED TIME = (OPTIMISTIC TIME + (4 * MOST LIKELY TIME) + PESSIMISTIC TIME) / 6

VARIANCE = ((PESSIMISTIC TIME - OPTIMISTIC TIME) / 6)^2


ACTIVITY

EXPECTED TIME

VARIANCE

A

(2 + (4 * 6) + 8) / 6 = 5.67

((8 - 2) / 6)^2 = 1

B

(10 + (4 * 11) + 12) / 6 = 11

((12 - 10) / 6)^2 = 0.11

C

(1 + (4 * 2) + 3) / 6 = 2

((3 - 1) / 6)^2 = 0.11

D

(6 + (4 * 7) + 8) / 6 = 7

((8 - 6) / 6)^2 = 0.11

E

(2 + (4 * 4) + 6) / 6 = 4

((6 - 2) / 6)^2 = 0.44

F

(1 + (4 * 2) + 3) / 6 = 2

((3 - 1) / 6)^2 = 0.11

G

(7 + (4 * 11) + 13) / 6 = 10.67

((13 - 7) / 6)^2 = 1

H

(11 + (4 * 12) + 13) / 6 = 12

((13 - 11) / 6)^2 = 0.11

I

(5 + (4 * 7) + 9) / 6 = 7

((9 - 5) / 6)^2 = 0.44

J

(14 + (4 * 16) + 18) / 6 = 16

((18 - 14) / 6)^2 = 0.44

K

(4 + (4 * 8) + 12) / 6 = 8

((12 - 4) / 6)^2 = 1.78